OpenStudy (anonymous):

x is less than or equal to 8 and greater than or equal to −5. Check my answer please!

OpenStudy (anonymous):

\[x \le8\ge-5\]

OpenStudy (steve816):

\[-5\le x \le 8\]

syed98 (syedmohammed98):

Agreed @steve816